OpenStudy (anonymous):

sin^3x+cos^3x over sinx+cosx = 1-sinx cosx

OpenStudy (lalaly):

\[\frac{\sin^3x+\cos^3x}{sinx+cosx}\] \[\sin^3x+\cos^3x=(sinx+cosx)(\sin^2x-sinx.cosx+\cos^2x)\] so\[=\frac{\cancel{(sinx+cosx)}(\sin^2x-sinxcosx+\cos^2x)}{\cancel{sinx+cosx}}\] \[ \sin^2x-sinxcosx+\cos^2x\] \[=1-sinxcosx\]